A Corporation is considering replacing the metal cutting machine (cutter) it currently uses to cut metal to make fences. The metal cutter has 6 years of remaining life. If kept, the metal cutter will have depreciation expenses of $800 for five years and $700 for the sixth year. Its current book value is $5,000 and it can be sold on eBay for $6,500 at this time. If the old metal cutter is not replaced it can be sold for $1,000 at the end of its useful life. Stuart Co. is considering purchasing the Fast Cutter XL, a higher-end metal cutter which costs $15,000 and has an estimated useful life of 6 years with an estimated salvage value of $2,000. This metal cutter falls into the MACRS 5-year class, so the applicable depreciation rates are 20.00%, 32.00%, 19,20%, 11.52%, 11.52% and 5.76%. The new cutter is faster and allows for an output expansion, so sales would rise by $2,500 per year, the new machine's much greater efficiency would reduce operating expenses by $1,500 per year. To support the greater sales, the new machine would require that inventories increase by $3,000 but accounts payable would simultaneously increase by $1,000. Stuart's marginal federal-plus-state tax rate is 25%, and its WACC is 12%.

What will be the after-tax salvage value cash flow of the existing (old) metal cutter if you sell it at the end of its natural life (6 years from now) and the after-tax salvage value cash flow of the existing (old) metal cutter if you sell it now?
Answer #1

Written down value after 6 years = Book value – Depreciation

= 5000-800*5 – 700

= $300

Sale value = $1000

Gain on sale = $700

Tax on gain = 700*25% = $175

After tax salvage value cash flow = Sale price – tax

= 1000-175

= $825

Gain if sold today = 6500-5000 = $1500

Tax on gain = 1500*25% = $375

After tax salvage value cash flow = 6500-375 = $6,125